# An effective $\boldsymbol{\Lambda}$-Szekeres modelling of the local Universe with Cosmicflows-4

### v1

@@ 1 Introduction
The effect of the presence of LSS on the large-scale background geometry and dynamics is commonly referred to as backreaction. It is conventionally quantified through the Buchert averaging scheme [Buchert_2000, Buchert_2001, Buchert_2006, Buchert_2018, Buchert_2020], under the working hypothesis that the FLRW metric well describes the averaged geometry on sufficiently large scales. In this context, backreaction appears as an effective modification of the Friedmann equations, introduced via additional source terms associated with the presence of nonlinear structures within the Universe [Buchert_2000, Buchert_2001, Buchert_2006, Buchert_2018, Buchert_2020]. Alternatively, mosaic universes—i.e., Swiss-Cheese models [Einstein_1945, Marra_2007, Biswas_2008], pancake solutions [Najera_2020], slab patchworks [Hellaby_2012], and cellular cosmologies [Lindquist_1957, Clifton_2012]—give a direct realisation of statistically homogenous universes which still display some LSS. Within these models, an emergent large-scale background geometry, in agreement with the Buchert averaging scheme predictions, is then recovered and can be directly compared to cosmological observations.
@@
At this present moment, no consensus has emerged within the scientific community on the effective impact of backreaction on cosmological observables [Clarkson_2011a, Kenworthy_2019, Clifton_2019]. Nonetheless, backreaction effects have been shown to correctly account for the dynamical behaviour of dark energy [Giani_2025a, Giani_2025b], and can even entirely replace the role of the latter in explaining the inferred acceleration of the Universe [Wiltshire_2007, Seifert_2024]. However, although a promising explanation for dynamical dark energy, backreaction effects from LSS cannot account for the anisotropies and small-scale inhomogeneities present in the local Universe.
@@
To model the impact of the local cosmic environment, we would ideally employ a model-independent framework such as the generalised cosmographic expansions developed in [Heinesen_2021, Kalbouneh_2024, Maartens_2024]. However, this approach suffers from a very limited radius of convergence in the presence of small-scale structures, i.e., $z_{\text{conv}}\ll 0.1$. Indeed, this limitation has been shown through investigations with general relativistic numerical simulations [Macpherson_2021, Macpherson_2023], Newtonian N-body simulation over an FLRW background [Adamek_2024], tests within the spherically symmetric Lemaître–Tolman–Bondi (LTB) spacetimes [Modan_2024], and Newtonian reconstructions of our local environment [Koksbang_2024]. Therefore, this scenario suggests that alternative approaches may be more suitable to assess the influence of the local Universe on our cosmological inferences.
@@
Another viable option is to directly model the local structures via exact solutions of the Einstein field equations (EFE). Such solutions must either naturally asymptote to an FLRW geometry on cosmological scales, or be intended to describe only the local environment and must be matched—possibly via a transition layer—to an FLRW model at larger scales [Krasiński_1997]. Over the past decade, this approach has attracted growing interest within the cosmological community, with a significant focus on spherically symmetric LTB solutions to model the impact of single structures, typically either a void or an overdensity, on cosmological observables [Camarena_2021, Camarena_2022, Marra_2022]. Moreover, anisotropic Szekeres solutions have been employed to describe small-scale single overdensity-void pairs [Bolejko_2008, Bolejko_2016], and a triaxial Bianchi I solution, inspired by a general anisotropic gravitational collapse [Giani_2021] has been implemented for an anisotropic modelling of the super-cluster Laniakea [Giani_2024a].
@@
In this paper, we adopt an analogous strategy, with the aim of obtaining an effective general-relativistic model of the local Universe as inferred from field-level reconstructions based on peculiar velocity (PV) data as tracers of the underlying matter distribution [Valade_2022, Courtois_2023, Valade_2024, Courtois_2025]. In particular, we consider the HAMLET reconstruction [Valade_2022, Valade_2024] based on the largest compilation of peculiar velocities currently available, namely the Cosmicflows-4 (CF4) catalogue [Tully_2023]. The local Universe presents high degrees of inhomogeneity and anisotropy, encoded in the presence of multiple structures. Here, we propose that such an intricate network of structures can be described, at first order, by a combination of multistructured $\Lambda$-Szekeres models [Krasiński_1997]. Furthermore, we assume that the large-scale geometry of the Universe is well approximated by an FLRW metric, and thus constrain our effective model by imposing FLRW boundary conditions. We fit multistructured Szekeres models to the peculiar PV and density HAMLET reconstruction of CF4, and extract the relevant dynamical variables. Using this effective description, we map the spatial curvature and expansion field within the local Universe. Additionally, we compute the impact of the inhomogeneous and anisotropic expansion rate within the local Universe on the redshift-distance relation for an observer located in the Milky Way. Finally, we use these corrections to investigate the impact of the local anisotropic expansion on cosmological observables.
@@
The remainder of this paper is structured as follows: in Sec. 2 we employ the CF4 velocity and density field reconstructions to fit multi-structured $\Lambda$-Szekeres models to the local Universe. In Sec. 3 we derive the corrections induced by the effective $\Lambda$-Szekeres model for the local structure on comoving distances, and their impact on cosmological observables, with a particular focus on the Hubble parameter, $H_{0}$. Sec. 4 is dedicated to the discussion of our findings and conclusions. Finally, in App. A we give a brief review of the formalism employed to describe the $\Lambda$-Szekeres solutions, and in App. B we present the code routine validation employed in this study.
@@ 2 An effective model for the local Universe
Our aim is to construct an effective description of the local Universe that qualitatively captures the impact of both anisotropies and inhomogeneities. To this end, we propose a toy model in which the cosmic environment is represented as a patchwork of inhomogeneous spherical wedges—akin to “pizza slices”. Within each wedge, local structures are modelled by a superposition of monopole and dipole components, allowing us to capture both the dominant isotropic contributions and leading-order directional deviations. With such a description, the local Universe behaves analogously to a class I $\Lambda$-Szekeres spacetime filled with dust and a cosmological constant [Szekeres_1975a, Szekeres_1975b, Sussman_2012, Sussman_2015, Sussman_2016, Sussman_2017, Buckley_2020]. In this framework, the dynamics of the local network of overdensities and voids—described within its rest frame—is driven by local differential expansion and collapse. The relative motion between any two points arises entirely from the inhomogeneous and anisotropic expansion field, i.e., a priori, there are no additional PVs sourced by local density gradients.
Furthermore, within our model we neglect the time evolution of both anisotropies and inhomogeneities, and instead model the local Universe by assuming that the background Hubble parameter is modified by a line-of-sight dependent function. This approximation is justified by the extremely low effective redshift of the map derived from the CF4 data set, where corrections due to time evolution can safely be neglected [Giani_2024a, Koksbang_2024]. Here, we consider radially-dependent line-of-sight corrections to the background Hubble parameter, which are necessary to capture the local Universe’s inhomogeneities. This is in contrast to the homogenous Laniakea model developed in Ref. [Giani_2024a], where local inhomogeneities were neglected.
@@ 2.1 Fields reconstruction from the Cosmicflows-4 data
The large-scale structure (LSS) formed by galaxies on scales of dozen to hundreds of Mpc has been first observed in the 1980s by the first systematic redshift surveys [deLapperent_etal_1986]. Since then, many more redshift surveys have extended in depth, spatial density and sky-covering our knowledge of the LSS [SDSS_Collaboration_2000, Jones_6dF_2009, DESICollaboration_2025]. Yet, while providing an intuitive picture of the so-called cosmic-web, this approach only maps the distribution of galaxies in the Universe, which are known to be biased tracers of the under-lying distribution of matter [Kaiser_1987]. Indeed, galaxies are found only at the peaks of the distribution of baryonic matter and thus do not directly trace dark matter, which is actually constituting the majority of the matter budget on these scales. Methods to reconstruct the density and velocity fields from pure redshift data exist and provide very convincing maps, e.g., [McAlpine_2025], but they necessarily rely on the assumption of a relationship between galaxy-number density and underlying total, gravitationally active matter density.
Considering the galaxies’ velocities instead of their positions presents a powerful advantage. Indeed, taken as tracers of the velocity field, galaxies are sensitive to the total distribution of gravitationally active matter, be it luminous or dark. Thus no galaxy-bias needs to be assumed. Yet, observations of peculiar velocities are much more challenging, leading to scarce, noisy data prone to interpretation biases [Strauss_1995, Nusser_2025]. the The Wiener-Filter / Constrained Realizations framework [Hoffman_Ribak_1992, Zaroubi_1999] has been the standard approach for the last decades, yet two inherent assumptions clash with the reality of the data, namely that (i) observational uncertainties are Gaussian, and (ii) that the positions of the constraints are known. This motivated the introduction of field-level forward modeling techniques, consisting of a Monte Carlo exploration of the posterior distribution of an arbitrary complex Bayesian model [Lavaux_2016, Graziani_2019, Valade_2022, Boruah_2022]. In practice, these models describe the fields in the same manner as the WF/CRs framework, but differ in their modeling of the observations.
These models—used to reconstruct the density and velocity fields—however still rely on two major assumptions. First, that the fields can be described as Gaussian and that the power spectrum of the density field is that of the CMB [Planck_2018]. Second, that the matter density perturbation field $\delta$ and the three-dimensional PV field $\vec{v}_{\mathrm{pec}}$ are related by the linearised continuity equation [KodamaSasaki_1984, Mukhanov_1992, Malik_2009]
@@
where $a(t)$ is the background scale factor. The density contrast field can be factorized as $\delta(\vec{x},t)=D_{+}(t)\delta_{0}(\vec{x})$, with $D_{+}(t)$ representing the linear growth factor of density fluctuations [KodamaSasaki_1984, Mukhanov_1992, Malik_2009]. Substituting this form into Eq. (2.1) leads to a convenient expression connecting the divergence of the PV to the density perturbation
@@
This assumption, commonly referred to as linear theory, is strictly valid only on large scales, of order tens of megaparsecs, although it may be extrapolated down to scales of a few megaparsecs at the cost of significantly increased uncertainties. In addition, applied outside its range of validity, linear theory may yield locally negative density in void regions and tends to over-estimate the contrast of structures in dense regions.
This work is based on density and velocity fields obtained from a field-level forward-modelling approach, namely the HAMLET reconstruction framework [Valade_2022, Valade_2024], applied to the largest compilation of peculiar-velocity measurements available to date, i.e., the CF4 catalogue. The dataset comprises more than 56,00056{,}000 peculiar velocities, providing nearly all-sky coverage out to redshift $z\simeq 0.05$, with particularly good sampling in the Galactic and celestial northern hemispheres extending to $z\simeq 0.1$ [Tully_2023]. The density and velocity fields are evaluated of a regular grid of $1\,{\rm Gpc/h}$ of side divided in 256 nodes, leading to a spacial resolution of $4\,{\rm Mpc/h}$.
@@ 2.2 Coarse-grained spatial geometry
In the $\Lambda$-Szekeres spacetimes, the dipole moment within the local matter distribution is encoded within via a single dipole function, $\boldsymbol{W}(r,\theta,\phi)$ (see App. A and, e.g., [Szekeres_1975b, Sussman_2012, Sussman_2015, Sussman_2016, Sussman_2017, Buckley_2020]). The dipole function is conventionally expressed as
@@
where X(r),Y(r),X(r),\,Y(r), and $Z(r)$ are arbitrary functions of the radial coordinate. We follow the ansatz of [Sussman_2015, Sussman_2016] for their functional form, i.e., in each $i$-th spherical wedge we take these functions to be piecewise defined over each $j$-th spherical shell by
@@
We emphasise that describing the local Universe through a spherical wedge-partition of multistructured $\Lambda$-Szekeres models can only, at best, serve as an effective approximation. This construction inherently introduces discontinuities at the interfaces between wedges—namely, layers of physical matter determined by the density gradient across the matching surfaces (see, e.g., [Sussman_2016]). However, we note that: (i) such discontinuities are already an unavoidable feature of any coarse-grained description of the local Universe; and (ii) in our approach, we neglect the time evolution of the nearby cosmic environment, so these artifacts ultimately play no role in our investigation. Indeed, the effective model we construct is intended only to probe how local spacetime expansion affects cosmological inferences. It is not meant to serve as a dynamical model of the local Universe’s evolution.
@@ 2.3 Effective $\boldsymbol{\Lambda}$-Szekeres reconstruction
The matter density $\rho$, the expansion rate scalar $H/3$, and the three-dimensional spatial curvature $\mathcal{K}$ in a $\Lambda$-Szekeres model are expressed within the quasilocal scalar formalism originally developed by Sussman and Bolejko [Sussman_2012], and further extended in Refs. [Sussman_2015, Sussman_2016, Sussman_2017], as
@@
where $\rho_{q}(t,r)$, $H_{q}(t,r)$, and $\mathcal{K}_{q}(t,r)$ denote proper volume averages—taken over spherical domains centred at the origin—of the corresponding physical scalars (see App. A and [Sussman_2012, Sussman_2015, Sussman_2016, Sussman_2017]). These quasilocal scalars then define a reference LTB background—able to capture the impact of even strong inhomogeneities—with respect to which the quantities $D^{(\rho)}$, $D^{(H)}$, and $D^{(\mathcal{K})}$ represent exact deviations that encode local anisotropies. The quasilocal scalars and their respective exact deviations satisfy the constraints [Sussman_2012, Sussman_2015, Sussman_2016, Sussman_2017]
@@
Therefore, from the PV data of CF4, we can directly extract the Hubble field, $H$, of the underlying $\Lambda$-Szekeres models, given an estimate of the expansion rate at the observer’s location. Here, we will consider as ansatz $\Delta H_{|r=0}=0$, i.e., we take the expansion within our cell to be, on average, the same as the fiducial cosmological models.
However, constructing a description in terms of quasilocal scalars and their exact deviations also requires determining the proper volume of the reconstructed cells, which is necessary for performing the weighted integration defining each quasilocal quantity. To determine the proper volume of the reconstructed cells, we exploit the linear nature of the employed CF4 reconstruction. Indeed, in this case we can find the proper volume, $\mathop{\text{d}\!}V_{\mathrm{prop}}$ as given by
@@
where $\mathop{\text{d}\!}V_{\mathrm{flat}}$ is the volume element in flat space, $\delta$ is the linear density contrast, and $a_{0}(r)$ is the current-time seed LTB scale-factor, which can be determined, given FLRW boundary conditions, to be [Sussman_2015, Sussman_2017]

### v2

@@ 1 Introduction
The effect of the presence of LSS on the large-scale background geometry and dynamics is commonly referred to as backreaction. It is conventionally quantified through the Buchert averaging scheme [12, 13, 8, 9, 10], under the working hypothesis that the FLRW metric well describes the averaged geometry on sufficiently large scales. In this context, backreaction appears as an effective modification of the Friedmann equations, introduced via additional source terms associated with the presence of nonlinear structures within the Universe [12, 13, 8, 9, 10]. Alternatively, mosaic universes — i.e., Swiss-Cheese models [39, 74, 4], pancake solutions [81], slab patchworks [51], and cellular cosmologies [68, 22] — give a direct realisation of statistically homogenous universes which still display some LSS. Within these models, an emergent large-scale background geometry, in agreement with the Buchert averaging scheme predictions, is then recovered and can be directly compared to cosmological observations.
@@
At this present moment, no consensus has emerged within the scientific community on the effective impact of backreaction on cosmological observables [21, 61, 23]. Nonetheless, backreaction effects have been shown to correctly account for the dynamical behaviour of dark energy [44, 46], and can even entirely replace the role of the latter in explaining the inferred acceleration of the Universe [117, 96]. However, although a promising explanation for dynamical dark energy, backreaction effects from LSS cannot account for the anisotropies and small-scale inhomogeneities present in the local Universe.
@@
To model the impact of the local cosmic environment, we would ideally employ a model-independent framework such as the generalised cosmographic expansions developed in [50, 60, 69]. However, this approach is complicated by the relatively small convergence radii of the cosmographic expansions, which are associated with the spatial gradients of small-scale inhomogeneities. This overall restricts the reliability of cosmographic distance estimates to low redshifts.Indeed, this limitation has been shown through investigations with general relativistic numerical simulations [70, 71], Newtonian N-body simulation over an FLRW background [1], tests within the spherically symmetric Lemaître–Tolman–Bondi (LTB) spacetimes [79], Newtonian reconstructions of our local environment [64], as well as in $\Lambda$-Szekeres models [52]. Although such an approach can still be employed to reliably study the presence of multipoles within the expansion field of the local Universe [59], alternative approaches may ultimately be better suited to assessing the influence of structures in the local Universe on our cosmological inferences.
@@
Another viable option is to directly model the local structures via exact solutions of the Einstein field equations. Such solutions must either naturally asymptote to an FLRW geometry on cosmological scales, or be intended to describe only the local environment and must be matched — possibly via a transition layer — to an FLRW model at larger scales [Krasiński_1997]. Over the past decade, this approach has attracted growing interest within the cosmological community, with a significant focus on spherically symmetric LTB solutions to model the impact of single structures, typically either a void or an overdensity, on cosmological observables [17, 18, 73, 16]. Moreover, anisotropic Szekeres solutions have been employed to describe small-scale single overdensity-void pairs [6, 5], and a triaxial Bianchi I solution, inspired by a general anisotropic gravitational collapse [45] has been implemented for an anisotropic modelling of the super-cluster Laniakea [43].
@@
In this paper, we adopt an analogous strategy, with the aim of obtaining an effective general-relativistic model of the local Universe as inferred from field-level reconstructions based on peculiar velocity (PV) data as tracers of the underlying matter distribution [112, 24, 111, 25]. In particular, we consider the HAMLET reconstruction [112, 111] based on the largest compilation of peculiar velocities currently available, namely the Cosmicflows-4 (CF4) catalogue [110]. The local Universe presents high degrees of inhomogeneity and anisotropy, encoded in the presence of multiple structures. Here, we propose that such an intricate network of structures can be described, at first order, by a combination of multistructured $\Lambda$-Szekeres models [Krasiński_1997]. Furthermore, we assume that the large-scale geometry of the Universe is well approximated by an FLRW metric, and thus constrain our effective model by imposing FLRW boundary conditions. We fit multistructured Szekeres models to the peculiar PV and density HAMLET reconstruction of CF4, and extract the relevant dynamical variables. Using this effective description, we map the spatial curvature and expansion field within the local Universe. Additionally, we compute the impact of the inhomogeneous and anisotropic expansion rate within the local Universe on the redshift-distance relation for an observer located in the Milky Way. Finally, we use these corrections to investigate the impact of the local anisotropic expansion on cosmological observables.
@@
The remainder of this paper is structured as follows: in Sec. 2 we employ the CF4 velocity and density field reconstructions to fit multi-structured $\Lambda$-Szekeres models to the local Universe. In Sec. 3 we derive the corrections induced by the effective $\Lambda$-Szekeres model for the local structure on comoving distances, and their impact on cosmological observables, with a particular focus on the Hubble parameter, $H_{0}$. Sec. 4 is dedicated to the discussion of our findings and conclusions. Finally, in App. A we give a brief review of the formalism employed to describe the $\Lambda$-Szekeres solutions, in App. B we show that the the coarse-graining procedure which we will employ in this work maintains the the large-scale structures within the CF4 input fields, and in App. C we present the code routine validation employed in this study.
@@ 2 An effective model for the local Universe
[…]
In this framework, the dynamics of the local network of overdensities and voids — described within its rest frame — is driven by local differential expansion and collapse. The relative motion between any two points arises entirely from the inhomogeneous and anisotropic expansion field, i.e., a priori, there are no additional PVs sourced by local density gradients. However, we stress that describing the reconstructed flow as peculiar velocities on top of an FLRW background, or equivalently as a local inhomogeneous expansion field, are two different but physically consistent descriptions of the same gravitationally-induced relative motions. Within the $\Lambda$-Szekeres interpretation, these motions are not removed; rather, they are encoded within the inhomogenous expansion field of an exact inhomogeneous spacetime. This then makes the framework particularly convenient for computing direction-dependent redshift–distance corrections within an internally consistent relativistic model, thus enabling a direct estimate of the impact on cosmological inference of the local cosmic neighbourhood.
Finally, we emphasise that within our model we neglect the time evolution of both anisotropies and inhomogeneities, and instead model the local Universe by assuming that the background Hubble parameter is modified by a line-of-sight dependent function. This approximation is justified by the extremely low effective redshift of the map derived from the CF4 data set, where corrections due to time evolution can safely be neglected [43, 64]. Here, we consider radially-dependent line-of-sight corrections to the background Hubble parameter, which are necessary to capture the local Universe’s inhomogeneities. This is in contrast to the homogenous Laniakea model developed in Ref. [43], where local inhomogeneities were neglected.
@@ 2.1 Fields reconstruction from the Cosmicflows-4 data
The large-scale structure (LSS) formed by galaxies on scales of dozen to hundreds of Mpc has been first observed in the 1980s by the first systematic redshift surveys [28]. Since then, many more redshift surveys have extended in depth, spatial density and sky-covering our knowledge of the LSS [94, 58, 33]. Yet, while providing an intuitive picture of the so-called cosmic-web, this approach only maps the distribution of galaxies in the Universe, which are known to be biased tracers of the underlying distribution of matter [35]. Indeed, galaxies are found only at the peaks of the distribution of baryonic matter and thus do not directly trace dark matter, which is actually constituting the majority of the matter budget on these scales. Methods to reconstruct the density and velocity fields from pure redshift data, such as the BORG implementation [57, 77] provide very convincing maps, but they necessarily rely on the assumption of a relationship between galaxy-number density and underlying total, gravitationally active matter density.
Considering the galaxies’ velocities instead of their positions presents a powerful advantage. Indeed, taken as tracers of the velocity field, galaxies are sensitive to the total distribution of gravitationally active matter, be it luminous or dark. Galaxies’ velocities have no bias at leading order. Yet, observations of peculiar velocities are relatively more challenging, leading to scarce, noisy data prone to interpretation biases [100, 82]. The Wiener-Filter / Constrained Realizations framework [54, 120] has been the standard approach for the last few decades, yet two inherent assumptions clash with the reality of the data, namely that (i) observational uncertainties are Gaussian, and (ii) that the positions of the constraints are known. This motivated the introduction of field-level forward modeling techniques, consisting of a Monte Carlo exploration of the posterior distribution of an arbitrary complex Bayesian model [67, 48, 112, 7]. In practice, these models describe the fields in the same manner as the WF/CRs framework, but differ in their modeling of the observations. Forward modeling methods, such as the one employed here, thus circumvent the use of ad-hoc velocity estimators such as [115, 53].
Linear reconstruction methods rely on two major assumptions. First, that the fields can be described as Gaussian and that the power spectrum of the density field is that of the CMB [87]. Second, that the matter density perturbation field $\delta$ and the three-dimensional PV field $\vec{v}_{\mathrm{pec}}$ are related by the linearized continuity equation [62, 80, 72]
@@
where $a(t)$ is the background scale factor. The density contrast field can be factorized as $\delta(\vec{x},t)=D_{+}(t)\delta_{0}(\vec{x})$, with $D_{+}(t)$ representing the linear growth factor of density fluctuations [62, 80, 72]. Substituting this form into Eq. (2.1) leads to a convenient expression connecting the divergence of the PV to the density perturbation
@@
This assumption, commonly referred to as linear theory, is strictly valid only on large scales, of order tens of megaparsecs, although it is often extrapolated down to scales of a few megaparsecs. Applied outside its range of validity, linear theory may yield locally negative density in void regions and tends to over-estimate the contrast of structures in dense regions.
[…]
Finally, the HAMLET reconstruction evaluates the density and velocity fields of a regular grid of $L_{\rm box}=1\,{\rm Gpc/h}$ side length divided into 256 nodes, leading to a spatial resolution of $4\,{\rm Mpc/h}$. Hence, the reconstructed volume is a priori sufficiently large enough to encompass the scales relevant to large-scale bulk flows [116, 115, 114], the major cosmic structures (i.e, walls, filaments and voids) within our cosmic neighbourhood (see e.g., [109, 88, 97, 119]), as well as various measures of homegeneity scale (for example [55, 84, 83, 36, 105, 42]). However, we note that the assumed periodicity of the box prevents the fitting of modes larger than the box-size, hence, some large scale power is lost. This effect is known to artificially dampen of $20\%$ (respectively $50\%$) the amplitude of the large scale bulk flow at $L_{\rm box}/10=100\,{\rm Mpc}/h$ (respectively $L_{\rm box}$[38]. Thus, we caution that the inferred results should only be compared to model predictions in the same volume. Finally, we note that within the HAMLET reconstruction $\Omega_{M}=0.3$, and $H_{0}=74.6~\mathrm{km\,s}^{-1}\,\mathrm{Mpc}^{-1}$ [111], i.e., the best fit cosmological parameters extracted for the CF4 data set [110]. Thus, no inconsistency bias is introduced at the level of background model parameter between data and reconstruction.
@@ 2.2 Coarse-grained spatial geometry
In the $\Lambda$-Szekeres spacetimes, the dipole moment within the local matter distribution is encoded within via a single dipole function, $\boldsymbol{W}(r,\theta,\phi)$ (see App. A and, e.g., [107, 101, 103, 102, 104, 15]). The dipole function is conventionally expressed as
@@
where X(r),Y(r),X(r),\,Y(r), and $Z(r)$ are arbitrary functions of the radial coordinate. We follow the ansatz of [103, 102] for their functional form, i.e., in each $i$-th spherical wedge we take these functions to be piecewise defined over each $j$-th spherical shell by
@@
We emphasise that describing the local Universe through a spherical wedge-partition of multistructured $\Lambda$-Szekeres models can only, at best, serve as an effective approximation. This construction inherently introduces discontinuities at the interfaces between wedges — namely, layers of physical matter determined by the density gradient across the matching surfaces (see, e.g., [102]). However, we note that: (i) such discontinuities are already an unavoidable feature of any coarse-grained description of the local Universe; and (ii) in our approach, we neglect the time evolution of the nearby cosmic environment, so these artifacts ultimately play no role in our investigation. Indeed, the effective model we construct is intended only to probe how local spacetime expansion affects cosmological inferences. It is not meant to serve as a dynamical model of the local Universe’s evolution.
@@ 2.3 Effective $\boldsymbol{\Lambda}$-Szekeres reconstruction
The matter density $\rho$, the expansion rate scalar $H/3$, and the three-dimensional spatial curvature $\mathcal{K}$ in a $\Lambda$-Szekeres model are expressed within the quasilocal scalar formalism originally developed by Sussman and Bolejko [101], and further extended in Refs. [103, 102, 104], as
@@
where $\rho_{q}(t,r)$, $H_{q}(t,r)$, and $\mathcal{K}_{q}(t,r)$ denote proper volume averages taken over spherical domains centred at the origin of the corresponding physical scalars (see App. A and [101, 103, 102, 104]). In this sense, the coordinate origin is thus interpreted as the observer-centred anchor of the effective reconstruction. Then, since the local Universe is inhomogeneous and anisotropic, shifting this origin (i.e., the observer in our framework) would in general lead to a different radial decomposition and hence to different quasilocal averages. Therefore, the present construction should be understood as an effective description of the cosmic neighbourhood as reconstructed from our observational position. These quasilocal scalars then define a reference LTB background — able to capture the impact of even strong inhomogeneities — with respect to which the quantities $D^{(\rho)}$, $D^{(H)}$, and $D^{(\mathcal{K})}$ represent exact deviations that encode local anisotropies. The quasilocal scalars and their respective exact deviations satisfy the constraints [101, 103, 102, 104]
@@
Therefore, from the PV data of CF4, we can directly extract the Hubble field, $H$, of the underlying $\Lambda$-Szekeres models, given an estimate of the expansion rate at the observer’s location. Here, we will consider as ansatz $\Delta H_{|r=0}=0$, i.e., we take the expansion within our cell to be, on average, the same as the fiducial cosmological models. This assumption — adopted consistently throughout the analysis — is motivated by the fact that the HAMLET reconstruction provides PVs in the comoving frame. As a result, the density and PV fields used to calibrate the effective $\Lambda$-Szekeres model are not affected by a dipole bias induced by the peculiar motion of the observer with respect to the Hubble flow.
Constructing a description in terms of quasilocal scalars and their exact deviations also requires determining the proper volume of the reconstructed cells, which is necessary for performing the weighted integration defining each quasilocal quantity. To determine the proper volume of the reconstructed cells, we exploit the linear nature of the employed CF4 reconstruction. Indeed, in this case we can find the proper volume, $\mathop{\text{d}\!}V_{\mathrm{prop}}$ as given by
@@
where $\mathop{\text{d}\!}V_{\mathrm{flat}}$ is the volume element in flat space, $\delta$ is the linear density contrast, and $a_{0}(r)$ is the current-time seed LTB scale-factor, which can be determined, given FLRW boundary conditions, to be [103, 104]